# Denoising as Adaptation: Noise-Space Domain Adaptation for Image Restoration

### v1

@@ Abstract
Although deep learning-based image restoration methods have made significant progress, they still struggle with limited generalization to real-world scenarios due to the substantial domain gap caused by training on synthetic data. Existing methods address this issue by improving data synthesis pipelines, estimating degradation kernels, employing deep internal learning, and performing domain adaptation and regularization. Previous domain adaptation methods have sought to bridge the domain gap by learning domain-invariant knowledge in either feature or pixel space. However, these techniques often struggle to extend to low-level vision tasks within a stable and compact framework. In this paper, we show that it is possible to perform domain adaptation via the noise-space using diffusion models. In particular, by leveraging the unique property of how the multi-step denoising process is influenced by auxiliary conditional inputs, we obtain meaningful gradients from noise prediction to gradually align the restored results of both synthetic and real-world data to a common clean distribution. We refer to this method as denoising as adaptation. To prevent shortcuts during training, we present useful techniques such as channel shuffling and residual-swapping contrastive learning. Experimental results on three classical image restoration tasks, namely denoising, deblurring, and deraining, demonstrate the effectiveness of the proposed method. Code will be released at: https://github.com/KangLiao929/Noise-DA/.
@@ 1 Introduction
Image restoration is a long-standing yet challenging problem in computer vision. It includes a variety of sub-tasks, e.g., denoising [1, 2, 3], deblurring [4, 5], and deraining [6, 7], each of which has received considerable research attention. Many existing methods are based on deep learning, typically following a supervised learning pipeline. Since annotated samples are not available in real-world contexts, i.e., degradation is unknown, a common technique is to generate synthetic low-quality data from high-quality images based on some assumptions on the degradation process to obtain training pairs. This technique has achieved considerable success but is not perfect, as synthetic data cannot cover all unknown or unpredictable degradation factors, which can vary wildly due to uncontrollable environmental conditions. Consequently, existing restoration methods often struggle to generalize well to real-world scenarios.
@@
Figure 1: (a) The prediction error of a diffusion model is highly dependent on the quality of the conditional inputs. In this experiment, we introduce an additional condition alongside the original noisy input. This condition is the same target image but corrupted with additive white Gaussian noise at a noise level $\sigma\in[0,80]$. More details can be found in the Appendix. (b) The restoration network is optimized to provide “good” conditions to minimize the diffusion model’s noise prediction error, aiming for a clean target distribution.
@@
Extensive studies have been conducted to address the lack of real-world training data. Some methods improve the data synthesis pipeline to generate more realistic degraded inputs for training [8, 9]. Other blind restoration approaches estimate the degradation kernel from the real degraded input during inference and use it as a conditional input to guide the restoration [10, 11]. Unsupervised methods [12, 13, 14, 15, 16, 5, 17] enhance input quality without relying on predefined pairs of clean and degraded images. These methods often use deep internal learning or self-supervised learning, where the model learns to predict clean images directly from the noisy or distorted data itself. In this paper, we investigate the problem assuming the existence of both synthetic data and real-world degraded images. This scenario fits a typical domain adaptation setting, where existing methods can be categorized into feature-space [18, 19, 20, 21, 22, 23] and pixel-space [24, 25, 26, 27] approaches. Both paradigms have their weaknesses: aligning high-level deep representations in feature space may overlook low-level variations essential for image restoration, while pixel-space approaches often involve computationally intensive adversarial paradigms that can lead to instability during training.
@@
In this work, we present a novel adaptation method for image restoration, which allows for a meaningful diffusion loss to mitigate the domain gap between synthetic and real-world degraded images. Our main idea stems from the observation shown in Fig. 1(a). Here, we measure the noise prediction error of a diffusion model conditioned on a noisy version of the target image. The trend in Fig. 1(a) shows that conditions with fewer corruption levels facilitate lower prediction errors of the diffusion model. In other words, “good” conditions give low diffusion loss, and “bad” conditions lead to high diffusion loss. While such a behavior may be expected, it reveals an interesting property of how conditional inputs could influence the prediction error of a diffusion model. Our method leverages this phenomenon by conditioning both the restored synthetic image and real image from a restoration network onto the diffusion model, as shown in Fig. 1(b). Both networks are jointly trained, with the restoration network optimized to provide “good” conditions to minimize the diffusion model’s noise prediction error, aiming for a clean target distribution. The goal of providing good conditions drives the restoration network to learn to improve the quality of its outputs. After training, the diffusion model is discarded, leaving only the trained restoration network for inference.
To bridge the gap between the restored synthetic and real outputs, our method carefully conceals the identity of the conditions. This prevents the diffusion model from simply learning to differentiate between synthetic and real conditions based on their channel index, avoiding a trivial shortcut in training. In addition, the pixel similarity between the noisy synthetic label and synthetic output is also easy to distinguish when they share the same clean image. To avoid the above shortcut learning, we design a channel shuffling layer at the beginning of the diffusion model. It randomly shuffles the channel index of synthetic and real-world conditions at each training iteration before concatenating them. We further propose a residual-swapping contrastive learning strategy to ensure the model genuinely learns to restore images accurately, rather than relying on easily distinguishable features.
@@
Our work represents the first attempt at addressing domain adaptation in the noise space for image restoration. We show the unique benefits offered by diffusion loss in eliminating the domain gap between the synthetic and real-world data, which cannot be achieved using existing losses. To verify the effectiveness of the proposed method, we conducted extensive experiments on three classical image restoration tasks, including denoising, deblurring, and deraining.
@@ 2 Related Work
Image Restoration. Image restoration aims to recover images degraded by factors like noise, blur, or data loss. Driven largely by the capabilities of various neural networks [28, 29, 30, 31, 32, 33], significant advancements have been made in sub-fields such as image denoising [34, 35, 36, 3, 15, 37], image deblurring [38, 39, 40, 41], and image deraining [42, 43, 44, 45]. In image restoration, loss functions are essential for training models. For example, the $L1$ loss minimizes average absolute pixel differences, ensuring pixel-wise accuracy. Perceptual loss uses pre-trained neural networks to compare high-level features, ensuring perceptual similarity. Adversarial loss involves a discriminator distinguishing between real and restored images, pushing the generator to create more realistic outputs. However, restoration models trained on synthetic images with these conventional loss functions still cannot escape from a significant drop in performance when applied to real-world domains.
@@
To address the mismatch between training and testing degradations, some supervised image restoration techniques [8, 9] improve the data synthesis pipeline, focusing on creating a training degradation distribution that balances accuracy and generalization in real-world scenarios. Some methods [10, 11] estimate and correct the degradation kernels to improve the restoration quality. Our work is orthogonal to these methods, aiming to bridge the gap between training and testing degradations.
@@
Unsupervised learning methods for image restoration leverage models that do not rely on paired training samples [12, 14, 15, 16, 5, 17, 46]. Techniques like Noise2Noise [12], Noise2Void [47], and Deep Image Prior [48] exploit the intrinsic properties of images, where the network learns to restore images by understanding the natural image statistics or by self-supervision. These unsupervised approaches have proven effective in restoration tasks, achieving impressive results comparable to supervised learning methods. However, they often struggle with handling highly complex or corrupted images due to their reliance on learned distributions and intrinsic image properties, which may not fully capture intricate details and show limited generalization to other restoration tasks.
@@
Domain Adaptation. The concept of domain adaptation is proposed to eliminate the discrepancy between the source domains and target domains [49, 50] to facilitate the generalization ability of learning models. Previous methods can be categorized into feature-space and pixel-space approaches. For example, feature-space adaptation methods [18, 19, 20, 21, 22, 23] adjust the extracted features from networks to align across different domains. Among these methods, some classical techniques are developed like minimizing the distance between feature spaces [18, 20] and introducing domain adversarial objectives [19, 21, 23]. Aligning high levels of deep representation may overlook crucial low-level variances that are essential for target tasks such as image restoration. In contrast, pixel-space domain adaptation methods [24, 25, 26, 27] achieve distribution alignment directly in the raw pixel level, by translating source data to match the “style" of a target domain. While they are easier to understand and verify for effectiveness from domain-shifted visualizations, pixel-space adaptation methods require careful tuning and can be unstable during training. Recent methods [51, 52, 53] compensate for the limitation of isolated domain adaptation by jointly aligning feature space and pixel space, shared the similar pipeline with CycleGAN [54]. However, they tend to be computationally demanding due to the need to train multiple networks (generally two generators and two discriminators) and the complexity of the cycle consistency loss. Different from the above feature-space and pixel-space methods, we propose a new noise-space solution that preserves low-level appearance across different domains with a compact and stable framework.
@@
Diffusion Model. Diffusion models [55, 56, 57] have gained significant attention as a novel approach in generative modeling. They work by gradually transforming a simple distribution (usually Gaussian) into a complex distribution in a series of steps, reversing the diffusion process. This approach shows remarkable success in text-to-image generation [58, 59, 60] and image restoration [61, 62, 63]. Often, conditions are fed to the diffusion model for conditional generation, such as text [58], class label [64], visual prompt [65], and low-resolution image [66], to facilitate the approximation of the target distribution. In this work, we show that the diffusion’s forward denoising process has the potential to serve as a proxy task to improve the model’s generalization ability in image restoration tasks.
Figure 2: Overview of the proposed framework. The restored synthetic and real-world images from the image restoration network are conditioned onto the diffusion model, adapting to the clean distribution in a multi-step denoising manner. Both networks are jointly trained, and the diffusion model is discarded after training. Gradients obtained from the diffusion model are used to drive the restoration network to produce better conditions.
@@ 3 Methodology
Problem Definition. We start by formulating the problem of noise-space domain adaptation in the context of image restoration. Given a labeled dataset [^1] Following the notations in domain adaptation, we use “label” to represent the ground truth image in the task of image restoration. from a synthetic domain and an unlabeled dataset from a real-world domain, we aim to train a model on both the synthetic and real data that can generalize well to the real-world domain. Supposed that ${\mathcal{D}}^{s}=\{({\bm{x}}_{i}^{s},{\bm{y}}_{i}^{s})\}_{i=1}^{N^{s}}$ denotes the labeled dataset containing $N^{s}$ samples from the source synthetic domain and ${\mathcal{D}}^{r}=\{{\bm{x}}_{i}^{r}\}_{i=1}^{N^{r}}$ denotes the unlabeled dataset with $N^{r}$ samples from the target real-world domain, where ${\bm{y}}^{s}$ is the clean image, ${\bm{x}}^{s}$ is the corresponding synthetic degraded image, and ${\bm{x}}^{r}$ is the real-world degraded image.
Image Restoration Baseline. The image restoration network can be generally formulated as a deep neural network $G(\cdot;{\bm{\theta}}_{G})$ with learnable parameter $\bm{\theta}_{G}$. This network is trained to predict the ground truth image ${\bm{y}}^{s}$ from its degraded observation $\bm{x}^{s}$ on the synthetic domain. The proposed noise space domain adaptation is not limited to a specific type of network architecture. One can choose from existing networks such as DnCNN [1], U-Net [67], RCAN [68], and SwinIR [33]. The approach is also orthogonal to existing loss functions used in image restoration, e.g., $L_{1}$ or $L_{2}$ loss, Charbonnier loss [69], perceptual loss [70, 71], and adversarial loss [30, 39]. To better validate the generality of the proposed approach, we adopt the widely used U-Net architecture and the Charbonnier loss, denoted as $\mathcal{L}_{Res}$, as our baseline. In the joint training, the diffusion model is trained using a diffusion objective, $\mathcal{L}_{Dif}$, while the restoration network is updated using both the $\mathcal{L}_{Res}$ and $\mathcal{L}_{Dif}$. The diffusion model is discarded after training.
@@ 3.1 Noise-Space Domain Adaptation
Ideally, the ground truth images and those restored images by an image restoration model from both synthetic and real-world data should lie in a shared distribution $\mathcal{S}$ of high-quality clean images. However, attaining such an ideal model that can universally map any degraded images onto the distribution $\mathcal{S}$, is exceedingly challenging. By assuming the restored images from synthetic and real-world data obey distinct distributions $\mathcal{S}^{s}$ and $\mathcal{S}^{r}$, our goal is to align $\mathcal{S}^{s}$ and $\mathcal{S}^{r}$ with $\mathcal{S}$. To this end, we introduce a diffusion model that conditions on $\mathcal{S}^{s}$ and $\mathcal{S}^{r}$ to approximate the target distribution $\mathcal{S}$. During training, this diffusion model is expected to guide the conditional distributions $\mathcal{S}^{s}$ and $\mathcal{S}^{r}$ toward $\mathcal{S}$.
@@
Given the commonly adopted case where the ground truth images from the synthetic dataset are available, we first explore adapting the target distribution $\mathcal{S}$ with a perspective of paired data. Without loss of generality, let us consider a synthetic degraded image $\bm{x}^{s}$ with its ground truth ${\bm{y}}^{s}$ from the synthetic domain and a real degraded image $\bm{x}^{r}$ from the real-world domain. Using the restoration network $G(\cdot;{\bm{\theta}}_{G})$, we can obtain the restored images ${\hat{\bm{y}}}^{s}$ and ${\hat{\bm{y}}}^{r}$, respectively. Then, we introduce a diffusion denoising process as a proxy task in the training process. It employs the predicted images ${\hat{\bm{y}}}^{s}$ and ${\hat{\bm{y}}}^{r}$ as conditions to help the diffusion model fit the distribution of $\mathcal{S}$. Following the notations in DDPM [56], we denote the diffusion model as $\bm{\epsilon}_{\theta}$ and formulate its optimization to the following objective:
@@
$\mathcal{L}_{Dif}=\mathbb{E}\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\left(\tilde{\bm{y}}^{s},\mathbf{C}(\hat{\bm{y}}^{s},\hat{\bm{y}}^{r}),t\right)\right\|_{2},$ (1)
@@
where $\tilde{\bm{y}}^{s}=\sqrt{\bar{\alpha}_{t}}{\bm{y}}^{s}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, $\bm{\epsilon}\sim N(0,\bm{I})$, $\bar{\alpha}_{t}$ is the hyper-parameter of the noise schedule, and $\mathbf{C}(\cdot,\cdot)$ denotes the concatenation operation along the channel dimension. During the joint training shown in Fig. 2, supervision from the diffusion loss in Eq. (1) will back-propagate to the conditions $\hat{\bm{y}}^{s}$ and $\hat{\bm{y}}^{r}$ if they are under-restored, i.e., far away from the expected distribution $\mathcal{S}$. This encourages the preceding restoration network to align $\hat{\bm{y}}^{s}$ and $\hat{\bm{y}}^{r}$ as closely as possible to $\mathcal{S}$.
@@
Figure 3: During joint training of image restoration network and diffusion network, the restored result of synthetic degraded images smoothly converges to the expected distribution over the epochs. However, the model tends to find shortcut learning on real-world images by matching the similarity between the conditions and the paired clean image or remembering the channel index. Consequently, the image restoration network learns to corrupt the high-frequency details in real-world images and the diffusion model learns to ignore them.
@@
The joint training, however, could lead to trivial solutions or shortcuts, as shown in Fig. 3. For example, it is easy to distinguish the “synthetic” and “real” conditions by the pixel similarity between $\hat{\bm{y}}^{s}$ and $\tilde{\bm{y}}^{s}$ or the channel index. Consequently, the restoration network will cheat the diffusion network by roughly degrading the high-frequency information in real-world images. As illustrated in Fig. 3(bottom), we identify three stages in this training process: (I) Diffusion network struggles to recognize which conditions aid denoising as both are heavily degraded, promoting the restoration network to enhance both; (II) Synthetic image is clearly restored and is easy to discriminate from its appearance; (III) The diffusion model distinguish between the conditions, leading the restoration network to focus on the synthetic data while ignoring the real-world data.

### v2

@@ Abstract
Although learning-based image restoration methods have made significant progress, they still struggle with limited generalization to real-world scenarios due to the substantial domain gap caused by training on synthetic data. Existing methods address this issue by improving data synthesis pipelines, estimating degradation kernels, employing deep internal learning, and performing domain adaptation and regularization. Previous domain adaptation methods have sought to bridge the domain gap by learning domain-invariant knowledge in either feature or pixel space. However, these techniques often struggle to extend to low-level vision tasks within a stable and compact framework. In this paper, we show that it is possible to perform domain adaptation via the noise space using diffusion models. In particular, by leveraging the unique property of how auxiliary conditional inputs influence the multi-step denoising process, we derive a meaningful diffusion loss that guides the restoration model in progressively aligning both restored synthetic and real-world outputs with a target clean distribution. We refer to this method as denoising as adaptation. To prevent shortcuts during joint training, we present crucial strategies such as channel-shuffling layer and residual-swapping contrastive learning in the diffusion model. They implicitly blur the boundaries between conditioned synthetic and real data and prevent the reliance of the model on easily distinguishable features. Experimental results on three classical image restoration tasks, namely denoising, deblurring, and deraining, demonstrate the effectiveness of the proposed method.
@@ 1 Introduction
Image restoration is a long-standing yet challenging problem in computer vision. It includes a variety of sub-tasks, e.g., denoising (Zhang et al., 2017; Yue et al., 2024), deblurring (Pan et al., 2016; Ren et al., 2020), and deraining (Fu et al., 2017; Wang et al., 2021), each of which has received research attention. Many methods are based on deep learning, typically following a supervised learning pipeline. Since annotated samples are not available in real-world contexts, i.e., degradation is unknown, a common technique is to generate synthetic low-quality data from high-quality images based on assumptions on the degradation process to obtain training pairs. This technique has achieved considerable success but is not perfect, as synthetic data cannot cover all unknown or unpredictable degradation factors, which can vary wildly due to uncontrollable environmental conditions. Consequently, existing methods often struggle to generalize well to real-world scenarios.
@@
Figure 1: (a) The prediction error of a diffusion model is highly dependent on the quality of the conditional inputs. In this experiment, we introduce an additional condition alongside the original noisy input. This condition is the same target image but corrupted with additive white Gaussian noise at a noise level $\sigma\in[0,80]$. More details can be found in the Appendix A1.1. (b) The restoration network is optimized to provide “good” conditions to minimize the diffusion model’s noise prediction error, aiming for a clean target distribution.
@@
Extensive studies have been conducted to address the lack of real-world training data. Some restoration methods improve the data synthesis pipeline to generate more realistic degraded inputs for training (Zhang et al., 2023; Luo et al., 2022). Other blind restoration approaches estimate the degradation kernel from the real degraded input during inference and use it as a conditional input to guide the restoration (Gu et al., 2019; Bell-Kligler et al., 2019). Unsupervised methods (Lehtinen et al., 2018; Shocher et al., 2018; Chen et al., 2023; Ren et al., 2020; Lee et al., 2022) enhance input quality without relying on predefined pairs of clean and degraded images. These methods often use deep internal learning or self-supervised learning, where the model learns to predict clean images directly from the noisy or distorted data itself. In this paper, we investigate the problem assuming the existence of both synthetic data and real-world degraded images. This scenario fits a typical domain adaptation setting, where existing methods can be categorized into feature-space (Tzeng et al., 2014; Ganin & Lempitsky, 2015; Long et al., 2015; Tzeng et al., 2015; Bousmalis et al., 2016) and pixel-space (Taigman et al., 2016; Shrivastava et al., 2017; Bousmalis et al., 2017) approaches. Both paradigms have their weaknesses: aligning high-level deep representations in feature space may overlook low-level variations essential for image restoration, while pixel-space approaches often involve computationally intensive adversarial paradigms that can lead to instability during training.
@@
In this work, we present a novel domain adaptation method for image restoration, which allows for a meaningful diffusion loss to mitigate the domain gap between synthetic and real-world degraded images. Our main idea stems from the observation shown in Fig. 1(a). Here, we measure the noise prediction error of a diffusion model conditioned on a noisy version of the target image. The trend in Fig. 1(a) shows that conditions with fewer corruption levels facilitate lower prediction errors of the diffusion model. In other words, “good” conditions give low diffusion loss, and “bad” conditions lead to high diffusion loss. While such a behavior may be expected, it reveals an interesting property of how conditional inputs could influence the prediction error of a diffusion model. Our method leverages this phenomenon by taming a diffusion model conditioned on both the restored synthetic image and restored real image from the restoration network, as shown in Fig. 1(b). Both networks are jointly trained, with the restoration network optimized to provide good conditions to minimize the diffusion model’s noise prediction error, aiming for a clean target distribution. Such a goal drives the restoration network to learn to improve the quality of its outputs. After training, the diffusion model is discarded, leaving only the trained restoration network for inference.
While the multi-step denoising process aids the restoration network, a potential shortcut learning could arise: the diffusion model learns to recognize conditions based on their channel index or pixel similarity to noisy synthetic labels, thereby neglecting real data. To mitigate this issue, we propose crucial strategies to fool the diffusion model, making it hard to discriminate between these two conditions. Specifically, we incorporate a channel-shuffling layer into the diffusion model and design a residual-swapping contrastive learning strategy to ensure the model genuinely learns to restore images accurately, rather than relying on easily distinguishable features. These strategies implicitly blur the boundaries between synthetic and real data, ensuring that both contribute effectively during joint training and facilitating their alignment with the target distribution.
To verify the effectiveness of our method, we conducted extensive experiments on three classical image restoration tasks (denoising, deblurring, and deraining), showing promising restoration performance and scalability to different networks. In summary, we make the following contributions:
@@
  • Our work represents the first attempt at addressing domain adaptation in the noise space for image restoration. We show the unique benefits from diffusion loss in eliminating the gap between the synthetic and real-world data, which cannot be achieved using existing losses.
  • To eliminate the shortcut learning in joint training, we design strategies to fool the diffusion model, making it difficult to distinguish between synthetic and real conditions, thereby encouraging both to align consistently with the target clean distribution.
  • Our method offers a general and flexible adaptation strategy applicable beyond specific restoration tasks. It requires no prior knowledge of noise distribution or degradation models and is compatible with various restoration networks. The diffusion model is discarded after training, incurring no extra computational cost during restoration inference.
@@ 2 Related Work
Image Restoration. Image restoration aims to recover images degraded by factors like noise, blur, or data loss. Driven largely by the capabilities of various networks (Dong et al., 2014; 2015; Zamir et al., 2022; Liang et al., 2021), significant advancements have been made in sub-fields such as image denoising (Zhang et al., 2021; Ren et al., 2021; Guo et al., 2019; Kim et al., 2020), image deblurring (Kupyn et al., 2018; Suin et al., 2020; Zhang et al., 2019), and image deraining (Jiang et al., 2020; Purohit et al., 2021; Ren et al., 2019; Yang et al., 2017). In image restoration, loss functions are essential for training models. For example, the $L1$ loss minimizes average absolute pixel differences, ensuring pixel-wise accuracy. Perceptual loss uses pre-trained networks to compare high-level features, ensuring perceptual similarity. Adversarial loss involves a discriminator distinguishing between real and synthetic images, pushing the generator to create more realistic outputs. However, restoration models trained on synthetic images with these conventional loss functions still cannot escape from a significant drop in performance when applied to real-world domains.
@@
To address the mismatch between training and testing degradations, some supervised image restoration techniques (Zhang et al., 2023; Luo et al., 2022) improve the data synthesis pipeline, focusing on creating a training degradation distribution that balances accuracy and generalization in real-world scenarios. Some methods (Gu et al., 2019; Bell-Kligler et al., 2019) estimate and correct the degradation kernels to improve the restoration quality. Our work is orthogonal to these methods, aiming to bridge the gap between training and testing degradations.
@@
Unsupervised learning methods for image restoration leverage models that do not rely on paired training samples (Huang et al., 2021; Chen et al., 2023; Huo et al., 2023; Chen et al., 2024). Techniques like Noise2Noise (Lehtinen et al., 2018), Noise2Void (Krull et al., 2019), and Deep Image Prior (Ulyanov et al., 2018) exploit the intrinsic properties of images, where the network learns to restore images by understanding the natural image statistics or by self-supervision. These approaches have proven effective in restoration tasks, achieving impressive results comparable to supervised learning methods. However, they often struggle with handling highly complex or corrupted images due to their reliance on learned distributions and intrinsic image properties, which may not fully capture intricate details and show limited generalization to other tasks.
@@
Domain Adaptation. The concept of domain adaptation is proposed to eliminate the discrepancy between the source domains and target domains (Saenko et al., 2010; Torralba & Efros, 2011) to facilitate the generalization ability of learning models. Previous methods can be categorized into feature-space and pixel-space approaches. For example, feature-space adaptation methods adjust the extracted features from networks to align across different domains. Among these methods, some classical techniques are developed like minimizing the distance between feature spaces (Tzeng et al., 2014; Long et al., 2015) and introducing domain adversarial objectives (Ganin & Lempitsky, 2015; Tzeng et al., 2017). Aligning high levels of deep representation may overlook crucial low-level variances that are essential for target tasks such as image restoration. In contrast, pixel-space adaptation methods (Liu & Tuzel, 2016; Taigman et al., 2016; Shrivastava et al., 2017; Bousmalis et al., 2017) achieve distribution alignment directly in the raw pixel level, by translating source data to match the “style" of a target domain. While they are easier to understand and verify for effectiveness from domain-shifted visualizations, pixel-space adaptation methods require careful tuning and can be unstable during training. Recent methods (Hoffman et al., 2018; Zheng et al., 2018; Chen et al., 2019) compensate for the limitation of isolated domain adaptation by jointly aligning feature space and pixel space. However, they tend to be computationally demanding due to the need to train multiple networks and the complexity of the cycle consistency loss (Zhu et al., 2017). Different from the above feature-space and pixel-space methods, we propose a new noise-space solution that preserves low-level appearance across different domains within a compact and stable framework.
@@
Diffusion Model. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Nichol & Dhariwal, 2021) have gained significant attention in generative modeling. They work by gradually transforming a simple distribution into a complex distribution in a series of steps, reversing the diffusion process. This approach shows remarkable success in text-to-image generation (Saharia et al., 2022b; Ruiz et al., 2023) and image restoration (Saharia et al., 2022a; c; Yue et al., 2023). Often, conditions are fed to the diffusion model for conditional generation, such as text (Rombach et al., 2021), class label (Ho & Salimans, 2022), visual prompt (Bar et al., 2022), and low-resolution image (Wang et al., 2023), to facilitate the approximation of the target distribution. Some recent works propose to adapt diffusion models for image restoration and its related tasks such as blind JPEG restoration (Welker et al., 2024), open-set image restoration (Gou et al., 2024), and classification of degraded images (Daultani et al., 2024). However, they require the diffusion model in both the training and inference stages. In this work, we show that the diffusion’s forward denoising process has the potential to serve as a training proxy task to improve the generalization ability of the image restoration model.
@@ 3 Methodology
Problem Definition. We start by formulating the problem of noise-space domain adaptation in the context of image restoration. Given a labeled dataset [^1] Following the notations in domain adaptation, we use “label” to represent the ground truth image. from a synthetic domain and an unlabeled dataset from a real-world domain, we aim to train a model on both the synthetic and real data that can generalize well to the real-world domain. Supposed that ${\mathcal{D}}^{s}=\{({\bm{x}}_{i}^{s},{\bm{y}}_{i}^{s})\}_{i=1}^{N^{s}}$ denotes the labeled dataset containing $N^{s}$ samples from the source synthetic domain and ${\mathcal{D}}^{r}=\{{\bm{x}}_{i}^{r}\}_{i=1}^{N^{r}}$ denotes the unlabeled dataset with $N^{r}$ samples from the target real-world domain, where ${\bm{y}}^{s}$ is the clean image, ${\bm{x}}^{s}$ is the corresponding synthetic degraded image, and ${\bm{x}}^{r}$ is the real-world degraded image.
Image Restoration Baseline. The image restoration network can be generally formulated as a deep neural network $G(\cdot;{\bm{\theta}}_{G})$ with learnable parameter $\bm{\theta}_{G}$. This network is trained to predict the ground truth image ${\bm{y}}^{s}$ from its degraded observation $\bm{x}^{s}$ on the synthetic domain. The proposed noise space domain adaptation is not limited to a specific type of network architecture. One can choose from existing networks such as DnCNN (Zhang et al., 2017), U-Net (Yue et al., 2019), RCAN (Zhang et al., 2018c), and SwinIR (Liang et al., 2021). The approach is also orthogonal to existing loss functions used in image restoration, e.g., $L_{1}$ or $L_{2}$ loss, Charbonnier loss (Zamir et al., 2021), perceptual loss (Johnson et al., 2016; Zhang et al., 2018b), and adversarial loss (Wang et al., 2018; Kupyn et al., 2018). To better validate the generality of the proposed approach, we adopt the widely used U-Net architecture and the Charbonnier loss, denoted as $\mathcal{L}_{Res}$, as our baseline. In the joint training, the diffusion model is trained using a diffusion objective, $\mathcal{L}_{Dif}$, while the restoration network is updated using both the $\mathcal{L}_{Res}$ and $\mathcal{L}_{Dif}$. The diffusion model is discarded after training.
@@ 3.1 Noise-Space Domain Adaptation
Ideally, the ground truth images and those restored images by an image restoration model from both synthetic and real-world data should lie in a shared distribution of high-quality clean images. However, attaining such an ideal model that can universally map any degraded images onto the distribution, is exceedingly challenging. Suppose a high-quality image $\bm{x}$ as a realization derives from a random vector $\bm{X}$, which belongs to the clean distribution $\bm{P_{X}}$. We then define the restored synthetic and real-world outputs from the restoration network as $\bm{\hat{X}^{s}}$ and $\bm{\hat{X}^{r}}$. In this work, we investigate developing a meaningful diffusion loss to guide the conditional distributions of both synthetic and real-world outputs aligned to the target clean distribution, i.e., $\bm{P_{X}}=\bm{P_{\hat{X}^{s}}}=\bm{P_{\hat{X}^{r}}}$.
@@
Given the commonly adopted case where the ground truth images from the synthetic dataset are available, we first explore adapting the target clean distribution with a perspective of paired data. Without loss of generality, let us consider a synthetic degraded image $\bm{x}^{s}$ with its ground truth ${\bm{y}}^{s}$ from the synthetic domain and a real degraded image $\bm{x}^{r}$ from the real-world domain. Using the restoration network $G(\cdot;{\bm{\theta}}_{G})$, we can obtain the restored images ${\hat{\bm{y}}}^{s}$ and ${\hat{\bm{y}}}^{r}$, respectively. Then, based on our observation that the predicted error of a diffusion model is highly dependent on the quality of the conditional inputs, we incorporate a multi-step denoising process as a proxy task into the training process. It employs the predicted images ${\hat{\bm{y}}}^{s}$ and ${\hat{\bm{y}}}^{r}$ as conditions to help the diffusion model fit the clean distribution. Following the notations in DDPM (Ho et al., 2020), we denote the diffusion model as $\bm{\epsilon}_{\theta}$ and formulate its optimization to the following objective:
@@
$\mathcal{L}_{Dif}=\mathbb{E}\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\left(\tilde{\bm{y}}^{s}|\mathbf{C}(\hat{\bm{y}}^{s},\hat{\bm{y}}^{r}),t\right)\right\|_{2},$ (1)
@@
where $\tilde{\bm{y}}^{s}=\sqrt{\bar{\alpha}_{t}}{\bm{y}}^{s}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, $\bm{\epsilon}\sim N(0,\bm{I})$, $\bar{\alpha}_{t}$ is the hyper-parameter of the noise schedule, and $\mathbf{C}(\cdot,\cdot)$ denotes the concatenation operation along the channel dimension. During the joint training, supervision from the diffusion loss in Eq. 1 will back-propagate to the conditions $\hat{\bm{y}}^{s}$ and $\hat{\bm{y}}^{r}$ if they are under-restored, i.e., far away from the expected distribution. This encourages the preceding restoration network to align $\hat{\bm{y}}^{s}$ and $\hat{\bm{y}}^{r}$ as closely as possible to the target clean distribution.
@@
Figure 2: During the joint training, the restored synthetic images smoothly converge to the expected distribution over the epochs. However, the model tends to find a shortcut in real data by matching the similarity between the conditions and the paired clean image or remembering the channel index. Consequently, the restoration network learns to corrupt the high-frequency details in real-world images and the diffusion model tends to ignore them.
@@
The joint training, however, could lead to trivial solutions or shortcuts, as shown in Fig. 2. For example, it is easy to distinguish the synthetic and real-world conditions by the pixel similarity between $\hat{\bm{y}}^{s}$ and $\tilde{\bm{y}}^{s}$ or the channel index. Consequently, the restoration network will cheat the diffusion network by roughly degrading the high-frequency information in real-world images. As illustrated in Fig. 2 (bottom), we identify three stages in this training process: (I) Diffusion network struggles to recognize which conditions aid denoising as both are heavily degraded, promoting the restoration network to enhance both; (II) Synthetic image is clearly restored and is easy to discriminate from its appearance; (III) The diffusion model tends to distinguish between the conditions, leading it to focus on the synthetic data while ignoring the real-world data.